OpenStudy (agent47):

@satellite73, same question as earlier part 2

OpenStudy (agent47):

P of heads = 1/3, total of 5 tosses, Z = X - Y. Compute E(Z) without computing the PMF of Z. I think I got it, just need the answer checked.. one second..

OpenStudy (agent47):

\[E(Z) = E(2X-5) = 2E(X)-E(5)=2*(1*1/3)-5=2/3-5\]that is

OpenStudy (anonymous):

i think there might be a mistake here

OpenStudy (agent47):

Hmm can you give me a hint, I want to try to understand what I'm doing wrong.

OpenStudy (anonymous):

ok just a hint how many heads do you expect if you flip 5 times ?

OpenStudy (agent47):

I'd expect 2.5, but we can't have that number, and the coin is biased so I'd lean towards 2 oh wait expected cant be negative lol

OpenStudy (anonymous):

it can for sure, for Z but that is not the question i was asking

OpenStudy (anonymous):

the fact that the coin is not fair should not throw you off

OpenStudy (agent47):

hang on i think i understand what you mean

OpenStudy (agent47):

E(X) = Sum over all possible x of x*(p(x)).. this will take me a while to do, I'm slow at prob

OpenStudy (agent47):

So that's the sum from 0 to 5

OpenStudy (anonymous):

oh no takes no time at all!!

OpenStudy (anonymous):

you are thinking way way too hard, these are "bernoulli" trials you flip five times you get heads one third of the time how many heads do you expect??

OpenStudy (agent47):

TTHTT so just one H?

OpenStudy (anonymous):

ok i think you are missing something here expected value does not have to be a possible outcome, it is just an average

OpenStudy (agent47):

Oh is it just the probability itself? Sorry I googled lol

OpenStudy (anonymous):

on average you get heads one third of the time you flip 300 times, how many heads would you expect?

OpenStudy (agent47):

I'd expect 100 heads

OpenStudy (anonymous):

right now suppose you only flip 100 times?

OpenStudy (agent47):

33.33333

OpenStudy (agent47):

5/3?

OpenStudy (agent47):

for 5 tosses

OpenStudy (anonymous):

right again

OpenStudy (agent47):

So \[E(Z)=E(2X-5)=2E(X)-E(5)=2*(5/3)-5=10/3-5=-5/3?\]

OpenStudy (anonymous):

kind of went off the page, but looks good

OpenStudy (agent47):

cool, thanks sat, I'll just write this down, there's also a part 3 for this thing, where I need to compute the expected value by computing the pmf first. Not sure how to approach the pmf yet. I'll tag you if you still have time.

OpenStudy (anonymous):

whatever \[\frac{10}{3}-5\] is

OpenStudy (anonymous):

you just do it

OpenStudy (anonymous):

the possible values of Z are what?

OpenStudy (agent47):

oh I see what you mean -5 <= z < = 5 and then calculate corresponding probabilities?

OpenStudy (anonymous):

right this approach, which will kinda suck, is meant for you to see how much easier the other is

OpenStudy (anonymous):

of course there are some values Z cannot take

OpenStudy (anonymous):

oh maybe just zero which is unimportant any way

OpenStudy (agent47):

Yea it can only be -5, -3, -1, 1, 3, 5

OpenStudy (anonymous):

ok looks like you are good to go

OpenStudy (agent47):

Ok, well I know how to finish this off, thanks sat. I have one last question a bit tougher, so I saved it for last. Can you take a look at that one please

OpenStudy (anonymous):

sure if i can what is it?